April 24, 2013 Tracking CDC/FDC
Meeting Time and Place
Wednesday April 24, 2013 at 11:00am At Jefferson Lab, the meeting will be held in F326
To connect from the outside, please use ESNET
- ) ESNET: 8542553
- )To connect by telephone, dial:
- You can look up toll-free number at http://www.readytalk.com/intl
- US and Canada: (866)740-1260 (toll free)
- International: (303)248-0285 (toll call)
- enter access code followed by the # sign: 3421244#
- FDC Update
- CDC Update CDC_odd_events
Present: Curtis, Naomi, Beni, Simon, Lubomir
- FDC Update
- Package #3 is almost ready to be moved to JLab for tests using the full electronics. Just missing one manifold that is being used in the cooling system tests.
- The set up for the test of the cooling system in the EEL126 is ready; still have to find a way to fill the chiller with Fluorinert and then do the test.
- Preparation of the mock-up to practice the cabling of the FDC at Blue Crab: half of the mesh cylinder is installed on one of the installation cards but may need some improvements.
- CDC Update
- Naomi reported on the CDC channel testing - so far 113 boards have been used for reading out data with HV of 2100 or 2200V, 50/50 Ar/CO2. Normally we would use 2.1kV but are going to 2.2kV now in order to save time later on if we should need to exceed 2.1kV.
- Only 2 sets of 3 transition boards have had trouble in reaching 2kV, all the others went to 2.2kV without a hitch. The trip setting used was 0.8uA total for 3 or 4 boards (72-96 straws). We will revisit the less cooperative boards and use a trip setting of 8uA if necessary. The boards usually draw very little current from the HV supply, 8nA or less for 3 boards is typical.
- Naomi has documented some unusual CDC events on the page CDC_odd_events. There are some more normal events shown on CDC_events. On page CDC_odd_events the top sets of events show pickup in strawless channels from huge hits on the straws, the pickup is about 1/10 x the amplitude of the hit signals; the lower sets of plots show some events where all 72 channels were subject to huge oscillations starting at the cosmic ray's arrival time. This was not seen on a subsequent run using the same straws. Apart from that one noisy run, this occurs in less than 1% of events.
- The daisy-chain LV supply cable for the preamps is starting to fall apart, 2 wires broke so far, Beni is going to look for some more cables and/or minibridge connectors.
- Lubomir will talk to Slava about finding drawings for the HV mounting blocks for the FDC in the hope that these can be copied for the CDC.
- fADC125 Update
- Beni heard from Cody that he has almost finished testing 30 fadcs - he still has to do the full crate tests - and hopes to be able to send some to CMU by the end of next week.